OpenStudy (anonymous):

what is the interest on $500 invested for 10 years in an account that earns simple interest at a rate of 2% per year ?

OpenStudy (imstuck):

Are you familiar with the formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

A = P(1 + rt) Where: A = Answer P = Principal R = Rate T = Time

OpenStudy (anonymous):

Plugin your numbers and you'll find the answer.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

500 = Principal Rate = 2% Time = 10 Years

OpenStudy (anonymous):

/so whats the answer

OpenStudy (anonymous):

Tell me what you get and I'll tell you if I got the same thing. We're not allowed to tell you.

OpenStudy (anonymous):

uhm idk how to solve it ???

OpenStudy (anonymous):

A = P(1 + rt) is a formula. You need to plug the information you're given into the formula and you'll get the answer. They give you the principal (p) the rate (r) and the time (t).

OpenStudy (anonymous):

your not getting what im saying do i divide or what

OpenStudy (anonymous):

What does the principal equal? What does the rate equal? What does the time equal?

OpenStudy (anonymous):

idkkkkkkk

OpenStudy (anonymous):

Did you read the question?

OpenStudy (anonymous):

more then once

OpenStudy (anonymous):

what is the interest on /$500\ invested for /10\ years in an account that earns simple interest at a rate of /2%\ per year ?

OpenStudy (imstuck):

I = prt

OpenStudy (anonymous):

What does the principal equal? What does the rate equal? What does the time equal?

OpenStudy (anonymous):

i just dont know if i divide first or what ?

OpenStudy (imstuck):

That's "Interest = Principle * Rate * Time

OpenStudy (imstuck):

Fill it in now.

OpenStudy (imstuck):

How much interest is what they want you to find.

OpenStudy (anonymous):

im confued give me the answer & ill work it out to find

OpenStudy (anonymous):

He's basically giving you the answer. Just plug the numbers in and do it.

OpenStudy (imstuck):

So, I = 500 * (the rate is the decimal form of 2%, which is .02) * 10

OpenStudy (anonymous):

I highlighted what p, r, and t equal. Now you just have to do basic math.

OpenStudy (anonymous):

im still confused do i divide ?

OpenStudy (anonymous):

* = multiply

OpenStudy (anonymous):

multiply 500mtimes 10?

OpenStudy (anonymous):

500 times 0.02

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

which is 10

OpenStudy (anonymous):

Okay now what is 10 times 10?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

There's your answer.

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

No problem :) hope I could help!

OpenStudy (anonymous):

can you help me one more time ? :)